5) Compute the magnitude and direction of the Coriolis force acting on a 1,000 Kg-car driving on I-15 towards Salt Lake at 80 mph.
Answer #1

Coriolis Force = 2 \Omega V M Sin\phi

\Omega is rotation rate of earth

V is the velocity ( 35.76 m/s)

2 \Omega = 1.458 * 10-4 s-1

M is the mass (1000 kg)

\phi is the latitude ( 40o)

Coriolis FOrce = 1.458 * 10-4 * 35.76 * 1000 * sin(40) = 3.88 Newton